Mr. McKelvey finds that flying with the wind he can travel 1188 miles in 6 hours. However, when flying against the wind, he travels only 2/3 of the distance is the same amount of time.
Find the speed of the plane in still air and the wind speed.

Answer by checkley71(8403)

1188/6=198 mph with the wind.
(1188*2/3)/6=792/6=132 mph against the wind
198-132=66 is the difference thus the wind is 66/2=33 mph.
proof 198-33=165 & 132+33=165 which is the speed of the plane in still air.